Q.1 Player 1 and player 2 bargain over sharing 400 dollars. The bargaining procedure follows the Rubinstein bargaining model. Player 1 makes the first offer. Player ls discount factor is 6, 1/2. Player 2s discount factor is 62-2/3. Find the bargaining solution
Answer #1

Answer:

Under Rubinstein Model

Player 1 will get x1 fraction and Player 2 will get 1-x2 fraction so

x1=(1-\delta 2)/(1-\delta 1*\delta 2)

So X1=(1-(1/2))/(1-(1/2)*(2/3))

X1=0.75

X2=1-0.75=0.25

So Player 1 will get X1*400=0.75*400=$300

So Player 2 will get X2*400=0.25*400=$100